OpenStudy (anonymous):

A board 60 in. long is cut into two parts so that the longer piece is 5 times the shorter. What are the lengths of the two pieces? The shorter piece is a0 in. The longer piece is a1 in.

OpenStudy (texaschic101):

l = longer piece s = shorter piece l + s = 60 l = 5s now sub 5s in for l in the other equation 5s + s = 60 6s = 60 s = 10 l = 5s l = 5(10) l = 50 shorter piece is 10 inches longer piece is 50 inches